OpenStudy (anonymous):

The temperature from a factory furnace varies inversely as the square of the distance from the furnace. The temperature 2 metres from the furnace is 50 degrees Celsius. Calculate the temperature 3.5 metres from the furnace. Give your answer to 2 decimal places.

OpenStudy (anonymous):

(1/d^2)K=T

OpenStudy (anonymous):

K(1/2^2)K=50 K(1/4)=50 K=200

OpenStudy (anonymous):

200(1/(3.5)^2)